OpenStudy (anonymous):

which complex number corresponds to 2e^ipi/6

OpenStudy (anonymous):

@sidsiddhartha

OpenStudy (sidsiddhartha):

use \[e^{i \theta}=\cos \theta+i \sin \theta\]

OpenStudy (anonymous):

cospi= -1 and sin pi= 0

OpenStudy (anonymous):

-1/3

OpenStudy (sidsiddhartha):

nope why ure taking pi , it will be pi/6 \[ \\ \huge2e^{i \frac{ \pi }{ 6 }}=2*[\cos (\pi/6)+i~\sin (\pi/6)]\]

OpenStudy (sidsiddhartha):

now simplify it

OpenStudy (sidsiddhartha):

got this?

OpenStudy (anonymous):

yes , 2[root 3/2]+isin[1/2]

OpenStudy (anonymous):

sorry 1/2i

OpenStudy (sidsiddhartha):

yup so it will be \[Z=\sqrt{3}+i\]

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

i have another question on vector lenghts and angles

OpenStudy (sidsiddhartha):

np :)

OpenStudy (sidsiddhartha):

go on

OpenStudy (anonymous):

i will have to attach a file, will do it in a min. Hang on kindly.

OpenStudy (anonymous):

question 4 pl

OpenStudy (sidsiddhartha):

but there is no number 4

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

i mean the three come under question 4

OpenStudy (anonymous):

i need the solutions for those

OpenStudy (sidsiddhartha):

consider any complex number-\[z=a+i~b ~~\\then ~~its~~vector~~lenght ~~is,R=\sqrt{a^2+b^2}\\and~~\angle ~~is,\theta=\tan^{-1} \frac{ b }{ a }\]

OpenStudy (sidsiddhartha):

just use this property

OpenStudy (anonymous):

so 3+ 4i, length is root(25)= 5

OpenStudy (anonymous):

angle=tan inverse (4/3)

OpenStudy (sidsiddhartha):

yup thats ok :)

OpenStudy (anonymous):

do we need to calculate arc tan ?

OpenStudy (sidsiddhartha):

yes u can calculate it

OpenStudy (anonymous):

for second part is it 6+0i?

OpenStudy (sidsiddhartha):

yes now here vector lenght,and angle is given , so again use eulers formula which is \[\large Z=r*e^{i \theta}=6*e^{i \frac{ \pi }{ 4 }}=6*[\cos(\pi/4)+i~\sin(\pi/4)]\]

OpenStudy (anonymous):

Thanks Sid

OpenStudy (sidsiddhartha):

no problem :)

OpenStudy (anonymous):

tomorrow by this time are u free ?

OpenStudy (anonymous):

i have doubts on Advance calculus 1

OpenStudy (sidsiddhartha):

sorry ireally cant say but i'll free from 7 pm to (whole night ) lol

OpenStudy (anonymous):

Thanks. let me try to hold u tomorrow. Thanks for todays session though.

OpenStudy (sidsiddhartha):

okey ^_^